OpenStudy (anonymous):

Simplify 2√2(-√8+2√6)

OpenStudy (anonymous):

I HAVE NO IDEA!!!!!!!!!

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

ansewer mine pleaseeeeeeee

OpenStudy (anonymous):

How do I see yours?

OpenStudy (solomonzelman):

\[2\sqrt{2}(-\sqrt{8}+2\sqrt{6})=-2\sqrt{16}+4\sqrt{12}=-2(4)+4(2)\sqrt{3}=-8+8\sqrt{3}\]

OpenStudy (anonymous):

\[2\sqrt{2}(2\sqrt{6}-2\sqrt{2)}\]

OpenStudy (anonymous):

That is not one of my options unfortuantly

OpenStudy (anonymous):

Oh sry...

OpenStudy (anonymous):

@solomon and alex

OpenStudy (anonymous):

Thanks though

OpenStudy (anonymous):

click my user name

OpenStudy (anonymous):

\[8\sqrt{3}-8\]

OpenStudy (anonymous):

maybe...

OpenStudy (anonymous):

give me ur options

OpenStudy (anonymous):

4√12-8 8-4√12 8√3-8

OpenStudy (anonymous):

oh so I was right :) Yay!

OpenStudy (anonymous):

Thank you! I'll let you know shortly if it really was right

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

was it right?